# Riemannian and Lorentzian Calderón problem under Magnetic Perturbation

### v1

@@ 1. Introduction
There exists a natural gauge equivalence for the DN map when $n\geq 3$. Namely, if $\Psi:M\to M$ is a diffeomorphism such that $\Psi|_{\partial M}$ is the identity map, then
@@
The well-known anisotropic Calderón problem asks if the metric $g$ is determined by the DN map $\Lambda_{g}$ up to boundary fixing isometry, we also refer to it as the Riemannian Calderón problem in this paper. The problem is widely open in dimension $n\geq 3$, see Section 1.1 for some history of the problem.
@@
Specifically, if $u$ is the solution with respect to $g$ and $A$, then $e^{i\omega}\Psi^{*}u$ is the solution with the same Dirichlet boundary data respect to $\Psi^{*}g$ and $\Psi^{*}A+d\omega$.
@@
Let $M$ be a smooth connected Riemannian manifold of dimension $n\geq 3$. Suppose $\Lambda_{g_{1},A}=\Lambda_{g_{2},A}$ for all $A$ in a neighborhood of 0 with respect to $C^{k}$ topology for some $k\geq 0$, then $g_{1}=g_{2}$.
The same problem can be considered when the metric is Lorentzian. Consider a Lorentzian manifold $(M,g)$ of dimension $n\geq 3$ with a timelike boundary $\partial M$. For simplicity, we assume the manifold is admissible (see Definition 3.1), in particular, the following wave equation is well-defined
@@
The wave equation (1.6) is well-posed for $f\in H^{1}_{loc}(\partial M)$ by [22, Theorem 24.1.1]. We define the DN map as
@@
where $\nu$ is the unit outward normal vector on the boundary, and $\bar{L}^{2}_{loc}(M^{\circ})$ refers to the space of restrictions to $M^{\circ}$ of locally $L^{2}$ functions on $M$ (see [22, Theorem 24.1.1] and [22, Page 113]). Again if $\Psi:M\to M$ is a boundary fixing diffeomorphism and $d\omega$ is a smooth exact 1-form such that $\omega|_{\partial M}=0$, then
@@
We refer to [40] for details of the gauge equivalences.
@@
If $L$ is the lens relation and $L(x^{\prime},\xi^{\prime})=(y^{\prime},\eta^{\prime})$, then the unique null-geodesic corresponding to $(y^{\prime},\eta^{\prime};x^{\prime},\xi^{\prime})$ is given by
@@
For the definition of principal symbol $\sigma(\Lambda_{g,A,q})$, see Section 3.1. A Lorentzian manifold is called non-trapping if any inextendible null-geodesic has two end points in the boundary. The assumption of the manifold being non-trapping is natural, since otherwise the information of the trapped region does not reach the boundary, and there is no way of telling the shape of light cone there. The admissibility assumption is to simplify the computation, and the idea of the proof should work for more general cases. Again we emphasize that the perturbation of $A$ eliminates the possibility of having boundary fixing isometry as a gauge equivalence.
Intuitively, the theorem states that perturbations of the electromagnetic potential in a test region can help determine whether the null-geodesic passes through it or not. We emphasize that the theorem is robust in the sense that if $x\in\gamma$, then a generic local perturbation of the electromagnetic potential in a neighborhood of $x$ would reveal this fact. Moreover, one does not need to have any a priori information about $\mathcal{A}$, the only thing required is the ability to add perturbation at various locations. It is worth mentioning that if $A-\mathcal{A}=d\omega$ and $\omega|_{\partial M}\equiv 0$, then by the gauge equivalence the DN map does not change. In fact the solution simply changes from $u$ to $Ce^{i\omega}u$ for some constant $C$. As mentioned before, in physics, this corresponds to $\mathcal{A}+d\omega$ and $\mathcal{A}$ represent the same electromagnetic field tensor $F$. On the other hand, they still represent the same electromagnetic field $F$ if $A-\mathcal{A}$ is a closed 1-form, but we no longer have the simple change in solutions as in the exact case. As a result, a change in closed 1-form may be used to detect the trajectory of null-geodesics. Specifically, in the proof we will show that the criterion is given by the shift in phase, in physics this is closely related to the well-known phenomenon called the Aharonov–Bohm effect, we refer to [1, 2].
@@
Suppose $(M,g)$ is non-trapping and admissible, denote $\Lambda_{g}$ the DN map of (1.11) with $q_{g}$ depending on $g$ locally. If $L$ is the lens relation and $L(x^{\prime},\xi^{\prime})=(y^{\prime},\eta^{\prime})$, then the unique null-geodesic associated with $(y^{\prime},\eta^{\prime};x^{\prime},\xi^{\prime})$ can be constructed by the following equivalence conditions.
@@ 1.1. History
The Calderón problem originated from the Electrical Impedance Tomography (EIT) problem considered by Calderón in [8]. EIT asks the following: given a bounded domain $\Omega\subset\mathbb{R}^{n}$ with an unknown (possibly anisotropic) conductivity tensor $\gamma(x)$, can one recover $\gamma$ from the Dirichlet-to-Neumann map which sends imposed voltages on $\partial\Omega$ to the resulting boundary currents on $\partial\Omega$? Equivalently, the DN map encodes exactly the voltage-to-current measurements at the boundary. In the anisotropic setting and for $n\geq 3$, this inverse boundary value problem is precisely Calderón’s original formulation of recovering the conductivity tensor from boundary data (up to the natural obstruction of boundary fixing diffeomorphisms). For the 2-dimensional Calderón problem, besides the boundary fixing isometry, a boundary fixing conformal change is also a gauge equivalence. The dimension 2 case has been fully solved in [29], but the higher dimension Calderón problem remains one of the central open questions in the theory of inverse problems. We refer to [38, 42] for more details on the history.
@@
When $n\geq 3$ and the metric is real analytic, the problem is solved in [28, 29, 31]. In [27], Lassas, Liimatainen and Salo proved the same result using a different approach. They used Poisson embedding to identify interior points with distributions on the boundary. When the metric is only $C^{\infty}$, in [11] and [12], it was shown that a metric in a fixed conformal class is uniquely determined by boundary measurements if the metric is conformally transversally anisotropic.
@@
The Lorentzian Calderón problem is relatively new comparing to the Riemannian version, but there has been some important progress in the past decade. One of the most important results is the recovery of the potential term in [3] and [4]. In [3], Alexakis, Feizmohammadi and Oksanen proved a type of unique continuation principle when the metric is given under some geometric assumptions, and used it to determine the potential term. When $A=0$ and the conformal class of $g$ is given, determining $g$ reduces to determining the potential term $q$. As a result, they recovered the conformal factor under some a priori conditions on the metric. In [4], they proved the same result for metrics sufficiently close to the Minkowski metric. For other results related to the recovery of potential term, see [39, 17, 24, 35]. Note that both of Theorem 1.2 and Theorem 1.3 recover the conformal class of the metric, one can then combine with results from [3] and [4] to recover the conformal factor.
@@
In [15, 16], Eskin proved the Lorentzian Calderón problem when the metric and potential are both real analytic in time. The main tool used there is Boundary Control method, which was introduced by Belishev in [5] and is closely related to the Unique Continuation Principle, see [7, 6, 41, 36] for details.
@@
As for the recovery of conformal class of the metric, one important result comes from when the equation has non-linearity. Specifically, in [25], Kurylev, Lassas and Uhlmann proved that the conformal class of the metric can be determined when the wave equation is non-linear and of the form
@@
They pioneered the use of higher order linearization and microlocal analysis to generate earliest light observation sets, which recovers the conformal class of the metric. The earliest light observation set is essentially the shape of the light cone intersected with the observation set. The method highly relies on the non-linearity, as it allows the waves to interact with each other and create a collision point in the interior. The same idea also helps recover the non-linearity itself, and the study of non-linear hyperbolic inverse problems has been an active research area ever since, for example see [45, 20, 9, 10, 33, 30, 46, 37, 44] and the surveys [26, 43]. This approach, however, does not work for linear equations, since two waves do not interact with each other.
Finally, there has been some boundary determination results. In [40], Stefanov and Yang proved that the local DN map is a pseudodifferential operator, which stably recovers the jet bundle of $g,A$ and $q$ on the boundary. They also proved that away from the diagonal, the DN map is a Fourier Integral Operator with canonical relation being the lens relation (see Section 3.1). Given the metric, they proved stable recovery of lens relation and light ray transform of $A$ and $q$ on the boundary. In [47], the author and Yang Zhang proved boundary determination under partial data assumption using the Fourier Integral Operator part of the DN map. Specifically, the Dirichlet initial data is supported on a fixed region $U$, while the observation is made on a fixed subset of the boundary $V$, and $U$ and $V$ are disjoint from each other. In both works, there is no interior recovery of the metric.
@@ 1.2. Idea of the proofs
We now briefly discuss the ideas behind the proofs and make comparison between the Riemannian and Lorentzian versions. As mentioned before, one of the key points in recovering the interior metric is the ability to focus on an interior point. Non-linearity and microlocal analysis in the Lorentzian setting provide this ability via wave interactions and propagation of singularity. However, in a linear equation the solutions do not interact, and in the Riemannian case there is no propagation.
@@
Suppose $B$ is an isomorphism on $T^{*}M$, and up to a diffeomorphism we may assume $B:T^{*}_{x}M\to T^{*}_{x}M$ for all $x\in M$. Suppose $B$ is the identity map on the boundary. Consider some Riemannian metric $g$, and suppose
$\Delta_{g}v=0\implies B(dv)\text{ is exact.}$. (1.14)
@@
Hence $B$ gives a gauge equivalence $\Lambda_{g}=\Lambda_{\widetilde{g}}$. We use Runge Approximation Theorem to show that the only such transformation in the Riemannian setting is the identity map, which proves Theorem 1.1.
@@ 2.1. Preliminaries
We first include a Runge Approximation Theorem result which will be used later, for a proof see for example the Appendix of [27].
Proposition 2.1 ([27] Proposition A.5).
@@
$-\Delta_{g}u=0\quad\text{in }M,\quad u|_{\partial M}=f$ (2.1)
@@ 2.2. Perturbation of the Magnetic Potential
We first recall the boundary determination result from [31]. The proposition has a specific construction formula, but we shall only state it as a deterministic result since it is all we need, for details we refer to the original paper.
Proposition 2.3 ([31] Proposition 1.3).
@@
[…]
@@
$\partial_{j}u=\frac{|g_{2}|^{\frac{1}{2}}}{|g_{1}|^{\frac{1}{2}}}(g_{1})_{jl}(g_{2})^{lk}\partial_{k}u_{2}:=B^{k}_{j}\partial_{k}u_{2}.$ (2.16)
@@
By Lemma 2.2, $B=CI$ for some constant $C\neq 0$ and $I$ is the identity map. Since $g_{1}$ and $g_{2}$ agree on the boundary by Proposition 2.3, this forces $C$ to be 1 and
$\frac{|g_{2}|^{\frac{1}{2}}}{|g_{1}|^{\frac{1}{2}}}g_{1}g_{2}^{-1}=I,$ (2.17)
meaning $g_{1}$ and $g_{2}$ are conformally equivalent. Take determinant we obtain
$|g_{1}|^{\frac{n-2}{2}}=|g_{2}|^{\frac{n-2}{2}},$ (2.18)
which gives $g_{1}=g_{2}$ since $n\geq 3$. ∎

### v2

@@ Abstract.
00footnotetext: Department of Mathematics, University of California San Diego. Email: yuyi@ucsd.edu
@@ 1. Introduction
Here $M^{\circ}$ is the interior of $M$. There exists a natural gauge equivalence for the DN map when $n\geq 3$. Namely, if $\Psi:M\to M$ is a diffeomorphism such that $\Psi|_{\partial M}$ is the identity map, then
@@
The well-known anisotropic Calderón problem asks if the metric $g$ is determined by the DN map $\Lambda_{g}$ up to boundary fixing isometry (see [LU89, Conjecture A]), we also refer to it as the Riemannian Calderón problem in this paper. The problem is widely open in dimension $n\geq 3$, see Section 1.1 for some history of the problem.
@@
This is referred to as the natural gauge equivalence. Specifically, if $u$ is the solution with respect to $g$ and $A$, then $e^{i\omega}\Psi^{*}u$ is the solution with the same Dirichlet boundary data respect to $\Psi^{*}g$ and $\Psi^{*}A+d\omega$.
@@
Let $M$ be a smooth connected compact manifold with boundary of dimension $n\geq 3$, and $g_{1},g_{2}$ are two Riemannian metrics on $M$. Suppose $\Lambda_{g_{1},A}=\Lambda_{g_{2},A}$ for all $A$ in a neighborhood of 0 with respect to $C^{k}$ topology for some $k\geq 0$, then $g_{1}=g_{2}$.
The following result is a direct consequence of Theorem 1.1.
Corollary 1.2.
For $j=1,2$, let $(M_{j},g_{j})$ be a smooth connected compact Riemannian manifold with boundary of dimension $n\geq 3$. Suppose there exists a boundary fixing diffeomorphism $\Psi:M_{1}\to M_{2}$ such that $\Lambda_{g_{1},A}=\Lambda_{g_{2},(\Psi^{-1})^{*}A}$ for all smooth 1-form $A$ on $M_{1}$ in a neighborhood of 0 with respect to $C^{k}$ topology for some $k\geq 0$, then $g_{1}=\Psi^{*}g_{2}$.
The same problem can be considered when the metric is Lorentzian. Consider a Lorentzian manifold $(M,g)$ of dimension $n\geq 3$ with a timelike boundary $\partial M$. We say the manifold is (lightlike) non-trapping if any inextendible lightlike geodesic $\gamma:[a,b]\to M$ eventually leaves $M$ in finite time, that is, $\infty<a\leq b<\infty$ and $\gamma(a),\gamma(b)\in\partial M$. For simplicity, we assume the manifold is non-trapping and admissible (see Definition 3.1), in particular, we consider the following electromagnetic wave equation
@@
The wave equation (1.6) is well-posed for $f\in H^{1}_{loc}(\partial M)$ by [Hor85III, Theorem 24.1.1]. We define the DN map as
@@
where $\nu$ is the unit outward normal vector on the boundary, and $\bar{L}^{2}_{loc}(M^{\circ})$ refers to the space of restrictions to $M^{\circ}$ of locally $L^{2}$ functions on $M$ (see [Hor85III, Theorem 24.1.1] and [Hor85III, Page 113]). Again if $\Psi:M\to M$ is a boundary fixing diffeomorphism and $d\omega$ is a smooth exact 1-form such that $\omega|_{\partial M}=0$, then
@@
We refer to [SY18] for details of the gauge equivalences.
@@
If $L$ is the lens relation (see Definition 3.2) and $L(x^{\prime},\xi^{\prime})=(y^{\prime},\eta^{\prime})$, then the unique null-geodesic corresponding to $(y^{\prime},\eta^{\prime};x^{\prime},\xi^{\prime})$ is given by
@@
Remark 1.4.
For the definition of principal symbol $\sigma(\Lambda_{g,A,q})$, see Section 3.1. A Lorentzian manifold is called non-trapping if any inextendible null-geodesic has two end points in the boundary. When the manifold is non-trapping and admissible (see Definition 3.1), all null-geodesics in the interior would eventually reach the boundary transversally, which makes the computation easier. We emphasize that the idea of the proof works for more general Lorentzian manifolds:
  1. (1)
    For example, the strictly convexity can be relaxed, then besides Proposition 3.3, one also needs to analyze the propagation of singularity along gliding rays or diffractive rays.
  2. (2)
    The non-trapping condition can also be removed, and one can only construct the null-geodesics that are non-trapping.
  3. (3)
    If the DN map is the partial data version as in [YZ25], then one can construct all the broken null-geodesics that travels from source region to observation region.
Finally, just like Theorem 1.1, since we work on a fixed base manifold, the perturbation of $A$ eliminates the possibility of having boundary fixing isometry as a gauge equivalence.
Intuitively, the theorem states that perturbations of the electromagnetic potential in a test region can help determine whether the null-geodesic passes through it or not. We emphasize that the theorem is robust in the sense that if $x\in\gamma$, then a generic local perturbation of the electromagnetic potential in a neighborhood of $x$ would reveal this fact. Moreover, one does not need to have any a priori information about $\mathcal{A}$, the only thing required is the ability to add perturbation at various locations. It is worth mentioning that if $A-\mathcal{A}=d\omega$ and $\omega|_{\partial M}\equiv 0$, then by the gauge equivalence the DN map does not change. In fact the solution simply changes from $u$ to $e^{i\omega}u$. In physics, this corresponds to $\mathcal{A}+d\omega$ and $\mathcal{A}$ represent the same electromagnetic field tensor $F$. On the other hand, they still represent the same electromagnetic field $F$ if $A-\mathcal{A}$ is a closed 1-form, but we no longer have the simple change in solutions as in the exact case. As a result, a change in closed 1-form may be used to detect the trajectory of null-geodesics. Specifically, in the proof we will show that the criterion is given by the shift in phase, in physics this is closely related to the well-known phenomenon called the Aharonov–Bohm effect, we refer to Remark 3.4 for more details. See also [AB59, AB61].
@@
Suppose $(M,g)$ is non-trapping and admissible, denote $\Lambda_{g}$ the DN map of (1.11) with $q_{g}$ depending on $g$ locally. If $L$ is the lens relation (see Definition 3.2) and $L(x^{\prime},\xi^{\prime})=(y^{\prime},\eta^{\prime})$, then the unique null-geodesic associated with $(y^{\prime},\eta^{\prime};x^{\prime},\xi^{\prime})$ can be constructed by the following equivalence conditions.
@@ 1.1. History
The Calderón problem originated from the Electrical Impedance Tomography (EIT) problem considered by Calderón in [Cal80]. EIT asks the following: given a bounded domain $\Omega\subset\mathbb{R}^{n}$ with an unknown (possibly anisotropic) conductivity tensor $\gamma(x)$, can one recover $\gamma$ from the Dirichlet-to-Neumann map which sends imposed voltages on $\partial\Omega$ to the resulting boundary currents on $\partial\Omega$? Equivalently, the DN map encodes exactly the voltage-to-current measurements at the boundary. In the anisotropic setting and for $n\geq 3$, this inverse boundary value problem is precisely Calderón’s original formulation of recovering the conductivity tensor from boundary data (up to the natural obstruction of boundary fixing diffeomorphisms). For the 2-dimensional Calderón problem, besides the boundary fixing isometry, a boundary fixing conformal change is also a gauge equivalence. The dimension 2 case has been fully solved in [LU01], but the higher dimension Calderón problem remains one of the central open questions in the theory of inverse problems. For more conductivity results, we refer to [SU87, KV84, KV85, HT13, AP06]; and we refer to [Sal13, Uhl14] for more details on the history.
@@
When $n\geq 3$ and the metric is real analytic, the problem is solved in [LTU03, LU01, LU89]. In [LLS20], Lassas, Liimatainen and Salo proved the same result using a different approach. They used Poisson embedding to identify interior points with distributions on the boundary. When the metric is only $C^{\infty}$, in [DKSU09] and [DKLS16], it was shown that a metric in a fixed conformal class is uniquely determined by boundary measurements if the metric is conformally transversally anisotropic.
@@
The Lorentzian Calderón problem is relatively new comparing to the Riemannian version, but there has been some important progress in the past decade. One of the most important results is the recovery of the potential term in [AFO22] and [AFO24]. In [AFO22], Alexakis, Feizmohammadi and Oksanen proved a type of unique continuation principle when the metric is given under some geometric assumptions, and used it to determine the potential term. When $A=0$ and the conformal class of $g$ is given, determining $g$ reduces to determining the potential term $q$. As a result, they recovered the conformal factor under some a priori conditions on the metric. In [AFO24], they proved the same result for metrics sufficiently close to the Minkowski metric. For other results related to the recovery of potential term, see [Ste89, FIKO21, Isa91, RS91]. Note that both of Theorem 1.3 and Theorem 1.5 recover the conformal class of the metric, one can then combine with results from [AFO22] and [AFO24] to recover the conformal factor.
@@
In [Esk07, Esk17], Eskin proved the Lorentzian Calderón problem when the metric and potential are both real analytic in time. The main tool used there is Boundary Control method, which was introduced by Belishev in [Bel87] and is closely related to the Unique Continuation Principle, see [BK92, Bel07, Tat99, RZ98] for details.
@@
As for the recovery of conformal class of the metric, one important result comes from when the equation has non-linearity. Specifically, in [KLU18], Kurylev, Lassas and Uhlmann proved that the conformal class of the metric can be determined when the wave equation is non-linear and of the form
@@
They pioneered the use of higher order linearization and microlocal analysis to generate earliest light observation sets, which recovers the conformal class of the metric. The earliest light observation set is essentially the shape of the light cone intersected with the observation set. The method highly relies on the non-linearity, as it allows the waves to interact with each other and create a collision point in the interior. The same idea also helps recover the non-linearity itself, and the study of non-linear hyperbolic inverse problems has been an active research area ever since, for example see [UZ23, HUZ22, CLOP22, CLOP21, OSSU24, LUW18, Yi24, SUW22, UZ22] and the surveys [Las25, UZ21]. This approach, however, does not work for linear equations, since two waves do not interact with each other.
Finally, there has been some boundary determination results. In [SY18], Stefanov and Yang proved that the local DN map is a pseudodifferential operator, which stably recovers the jet bundle of $g,A$ and $q$ on the boundary. They also proved that away from the diagonal, the DN map is a Fourier Integral Operator with canonical relation being the lens relation (see Section 3.1). Given the metric, they proved stable recovery of lens relation and light ray transform of $A$ and $q$ on the boundary. In [YZ25], the author and Zhang proved boundary determination under partial data assumption using the Fourier Integral Operator part of the DN map. Specifically, the Dirichlet initial data is supported on a fixed region $U$, while the observation is made on a fixed subset of the boundary $V$, and $U$ and $V$ are disjoint from each other. In both [SY18] and [YZ25], there is no interior recovery of the metric.
To the author’s knowledge, metric recovery from a family of DN maps indexed by the perturbation of parameters appears to be new in both Riemannian and Lorentzian settings. Classical Calderón results typically assume a single DN map and focus on boundary determination or single-map interior uniqueness. Almost all of the existing results are obtained under additional structural assumptions of the metric (for example, analyticity [LU01], CTA structures [DKSU09, DKLS16], or curvature bounds [FIKO21]), and many of them require the metrics to be in the same conformal class. In contrast, we add no such assumptions on the metric structure, but instead work with an active, coefficient-parametric data model: an open family of magnetic or electromagnetic perturbations and the associated family of DN maps.
@@ 1.2. Idea of the proofs
We now briefly discuss the ideas behind the proofs and make comparison between the Riemannian and Lorentzian versions. As mentioned before, one of the key points in recovering the interior metric is the ability to focus on an interior point. Non-linearity and microlocal analysis in the Lorentzian setting provide this ability via wave interactions and propagation of singularities. However, in a linear equation the solutions do not interact, and in the Riemannian case there is no propagation.
@@
Suppose $B$ is an isomorphism on $T^{*}M$, and up to a diffeomorphism we may assume $B:T^{*}_{x}M\to T^{*}_{x}M$ for all $x\in M$. Suppose $B$ is the identity map on the boundary, and the boundary is connected. Consider some Riemannian metric $g$, and suppose
$\Delta_{g}v=0\implies B(dv)\text{ is exact.}$ (1.14)
@@
Hence $B$ gives a gauge equivalence $\Lambda_{g}=\Lambda_{\widetilde{g}}$. We use a Runge Approximation type of result to show that the only such transformation in the Riemannian setting is the identity map, which proves Theorem 1.1.
@@ 2.1. Preliminaries
We first include a Runge Approximation type of result which will be used later, for a proof see for example the Appendix of [LLS20].
Proposition 2.1 ([LLS20, Proposition A.5]).
@@
$\Delta_{g}u=0\quad\text{in }M,\quad u|_{\partial M}=f$ (2.1)
@@ 2.2. Perturbation of the Magnetic Potential
We first recall the boundary determination result from [LU89]. The proposition has a specific construction formula, but we shall only state it as a deterministic result since it is all we need, for details we refer to the original paper.
Proposition 2.3 ([LU89, Proposition 1.3]).
@@
$\int_{M}\varphi^{\delta}\omega(\nabla_{g}u)|g|_{x}^{\frac{1}{2}}dx\xrightarrow{\delta\to 0}|g(x_{0})|^{\frac{1}{2}}\omega(\nabla_{g}u)|_{x_{0}},$ (2.11)
@@
$\partial_{j}u_{1}=\frac{|g_{2}|^{\frac{1}{2}}}{|g_{1}|^{\frac{1}{2}}}(g_{1})_{jl}(g_{2})^{lk}\partial_{k}u_{2}:=B^{k}_{j}\partial_{k}u_{2}.$ (2.16)
@@
By Lemma 2.2, $B=CI$ for some constant $C\neq 0$ and $I$ is the identity map. That is,
$\frac{|g_{2}|^{\frac{1}{2}}}{|g_{1}|^{\frac{1}{2}}}g_{1}g_{2}^{-1}=CI.$ (2.17)
Since we assumed the dimension $n\geq 3$, take determinant both sides and simplify, we obtain
$\frac{|g_{2}|^{\frac{1}{2}}}{|g_{1}|^{\frac{1}{2}}}=C^{\frac{n}{n-2}}.$ (2.18)
Hence
$g_{1}g_{2}^{-1}=C^{-\frac{2}{n-2}}I.$ (2.19)
Use the fact that $g_{1}$ and $g_{2}$ agree on boundary tangential directions by Proposition 2.3, $C=1$. That is, $g_{1}=g_{2}$. ∎